# Representation Meets Optimization: Training PINNs and PIKANs for Gray-Box Discovery in Systems Pharmacology

### v1

@@ Abstract
Physics-Informed Kolmogorov–Arnold Networks (PIKANs) are gaining attention as an effective counterpart to the original multilayer perceptron-based Physics-Informed Neural Networks (PINNs). Both representation models can address inverse problems and facilitate gray-box system identification. However, a comprehensive understanding of their performance in terms of accuracy and speed remains underexplored. In particular, we introduce a modified PIKAN architecture, tanh-cPIKAN, which is based on Chebyshev polynomials for parametrization of the univariate functions with an extra nonlinearity for enhanced performance. We then present a systematic investigation of how choices of the optimizer, representation, and training configuration influence the performance of PINNs and PIKANs in the context of systems pharmacology modeling. We benchmark a wide range of first-order, second-order, and hybrid optimizers—including various learning rate schedulers. We use the new Optax library [1] to identify the most effective combinations for learning gray-boxes under ill-posed, non-unique, and data-sparse conditions. We examine the influence of model architecture (MLP vs. KAN), numerical precision (single vs. double), the need for warm-up phases for second-order methods, and sensitivity to the initial learning rate. We also assess the optimizer scalability for larger models and analyze the trade-offs introduced by JAX in terms of computational efficiency and numerical accuracy. Using two representative systems pharmacology case studies — a pharmacokinetics model and a chemotherapy drug-response model — we offer practical guidance on selecting optimizers and representation models/architectures for robust and efficient gray-box discovery. Our findings provide actionable insights for improving the training of physics-informed networks in biomedical applications and beyond.
@@
PINNs, PIKANs, Kolmogorov-Arnold Networks, Systems Biology, Systems Pharmacology, Pharmacokinetics, QSP
@@ 1 Introduction
Inverse problems play a central role in scientific research, allowing practitioners to deduce hidden parameters and system behaviors based on empirical or experimental observations. These problems are of particular importance in fields such as biology, physics, and pharmacology, where direct measurement is often impractical and the underlying systems exhibit intricate nonlinear dynamics. Traditional methods like Sparse Identification of Nonlinear Dynamics (SINDy) [2], which rely on techniques from sparse regression [3] and compressed sensing [4], have been widely used to extract interpretable equations from data. However, these approaches often depend on predefined libraries of candidate functions and prior structural assumptions about the system. Addressing these limitations, the recently proposed ADAM-SINDy framework introduces a global optimization approach that simultaneously tunes nonlinear parameters and coefficients using the ADAM optimizer [5], significantly improving the robustness and adaptability of the standard SINDy method in handling complex dynamics. Methods like AI-Feynman [6], Feyn [7, 8], and AI-Descartes [9] focus on discovering symbolic expressions directly from data, typically without relying on predefined physical models. These are particularly useful in low-data regimes but face limitations when dealing with systems characterized by derivatives or integral terms. More recently, hybrid modeling approaches have emerged to combine the strengths of data-driven learning with physics-based constraints. Frameworks like PINNs [10] and their variants (e.g., X-PINNs [11]) have been integrated with symbolic regression to recover gray-box components in differential equation models [12]. Other notable contributions include the use of random projection neural networks (RPNNs) and interaction transformations [13, 14], which reduce computational complexity while retaining interpretability, as well as efficient symbolic-layer methods based on least-squares optimization [15].
Pioneering work on gray-box modeling using neural networks was introduced in [16], which highlighted the advantages of continuous-time system modeling for capturing complex nonlinear behaviors. These ideas have since been extended to diverse application areas, including phase-field systems, optogenetics, and biotechnology [17, 18, 19, 20, 21]. In biological systems, the challenge of deriving parameters from sparse and noisy data has been met through the employment of systems-biology-informed deep learning methodologies [22, 23]. This approach integrates the interpretability characteristic of mechanistic models with the flexibility inherent in neural networks, thereby addressing the limitations of traditional approaches. Most recently, a hybrid framework called AI-Aristotle was proposed [24] to integrate domain-decomposed PINNs and X-TFC [25] with symbolic regression to perform gray-box identification and parameter estimation in systems biology. Tested on pharmacokinetics and endocrine modeling tasks, AI-Aristotle demonstrated strong accuracy, robustness, and interpretability — even under sparse and noisy data regimes — highlighting the potential of combining physics-based machine learning with symbolic tools for complex dynamical system discovery. Among the pioneering efforts to apply physics-informed machine learning to system pharmacology modeling, compartment model-informed neural networks (CMINNs) [26] represent a significant advancement in solving inverse problems, where the parameters the parameters the parameters the parameters pharmacokinetic and pharmacodynamic (PK/PD) parameters are not constant. Unlike conventional compartmental models that assume fixed parameters, CMINNs integrate PINNs and fractional PINNs (fPINNs) with ordinary and fractional-order differential equations to capture complex drug behaviors such as anomalous diffusion, delayed effects, drug resistance, and persistence. By allowing parameters to be time-varying, constant, or piecewise constant — and even modeling the fractional derivative order as a learnable quantity — this framework enables more accurate and explainable modeling of drug absorption and response, particularly under multi-dose conditions in cancer therapy. This approach not only broadens the interpretability of model outputs but also enhances our understanding of non-standard drug kinetics in heterogeneous biological environments. Alongside AI-Aristotle, which focuses on gray-box identification using a hybrid of neural networks and symbolic regression, CMINNs stand as a foundational contribution to the field of inverse problem-solving in systems pharmacology—each tackling the challenge under different assumptions and modeling settings.
A major obstacle in neural network optimization is the non-convexity of the problem, which complicates the search for global minima. As a result, significant research has focused on understanding the training dynamics of neural networks [27, 28, 29]. Optimization methods are typically divided into three categories: first-order techniques, such as stochastic gradient descent [30, 31], high-order techniques, like Newton’s method [32, 33, 34], and heuristic, derivative-free approaches [35, 36]. Although high-order methods can improve convergence speed by utilizing curvature information, they encounter difficulties related to the computational overhead of calculating and storing the inverse of the Hessian matrix. To mitigate this issue, variations of Newton’s method, including BFGS and L-BFGS (Limited-memory BFGS), use rank-one or rank-two updates from prior gradient information to approximate the Hessian or its inverse [37, 38, 39, 40, 41, 42, 43].
Physics-Informed Neural Networks (PINNs) [44] have significantly advanced the solution of both inverse and forward problems in differential equations, particularly for systems exhibiting complex nonlinear behavior, numerous unknown parameters, and limited experimental data [45, 46]. The training process involves minimizing a composite loss function that simultaneously enforces the governing physical laws and aligns predictions with observed data, including constraints such as boundary, initial conditions, and, where available, observational data. The incorporation of a physics loss term in the optimization process introduces additional complexities. Recent studies have explored the impact of optimization algorithms in addressing these challenges. For example, research has shown that the application of advanced optimization algorithms can significantly improve the accuracy of the outputs obtained by both PINNs and PIKANs when solving forward problems [47, 48]. Training PINNs presents notable challenges across a wide range of scientific and engineering applications. One of the main difficulties stems from the inclusion of physics-based residual loss terms, which must be minimized alongside data-driven objectives. This creates a complex optimization landscape with multiple competing loss components, often leading to convergence issues, slow training, and the risk of getting trapped in local minima. These challenges become even more pronounced in systems pharmacology and systems biology, where inverse problems involve sparse, noisy, and often unbalanced datasets, and where unknown parameters may be time-varying or embedded within nonlinear dynamics. In AI-Aristotle [24], CMINNs [26] and [49] we addressed these issues by introducing a robust training framework tailored to these domains. Key strategies included a two-step training procedure, which begins with a warm-up phase focused on fitting the data before incorporating the full physics-informed loss, as well as the use of adaptive loss weighting, exponential feature layers, scaling layers, and sequential training. These techniques collectively improved convergence, reduced sensitivity to local minima, and enhanced model generalization without requiring changes to the network architecture or optimization algorithm. These solutions reflect broader trends and challenges in the PINNs literature, as also discussed in recent reviews [50, 45].
However, we observed that using first-order optimizers like Adam, with a constant learning rate, led to the model converging very slowly and often times getting stuck in local minima with convergence stalled. Even with extended training, the accuracy of the solution of the inverse problem could not be improved. Therefore, in this study, we aim to evaluate the impact of adaptive learning rates (based on different types of schedulers) for first-order optimizers and compare them with second-order optimizers such as BFGS with two line search algorithms backtracking, and trust-region. Additionally, we also investigate the effect of float (32 Bits) and double precision (64 Bits) arithmatic on the convergence of optimizers to the solution.
@@
The paper is organized as follows. In Section 2, we present the methodology, detailing the architectures of PINNs and the proposed tanh-cPIKANs, along with their respective formulations and training strategies. Section 3 introduces the pharmacokinetics and pharmacodynamics models used for evaluation, both of which represent gray-box inverse problems characterized by data sparsity, non-uniqueness, and ill-posedness. Section 4 provides a systematic investigation into representation and optimization choices for gray-box discovery. It compares the performance of MLP-based PINNs and KAN-based architectures, explores the influence of numerical precision (single vs. double) on convergence and stability, and evaluates the effectiveness of various optimizers and learning rate schedulers. The section also analyzes the necessity of warm-up phases for second-order methods, the sensitivity to initial learning rates, and the scalability of different optimization strategies. Additionally, it assesses the computational cost and accuracy trade-offs associated with using the JAX framework. Finally, Section 5 summarizes the key findings, emphasizing the advantages of the proposed tanh-cPIKANs architecture in enhancing training stability and accuracy for gray-box system identification.
@@ 2 Methodology
Physics-Informed Networks (PINs) are models that embed physical laws—typically represented by ordinary or partial differential equations (ODEs/PDEs)—directly into the training process of function approximators such as neural networks or Kolmogorov–Arnold Networks. This is accomplished through automatic differentiation, which enables the computation of derivatives required to enforce consistency with the governing equations. The term “PINs” broadly encompasses both Physics-Informed Neural Networks (PINNs) and Physics-Informed Kolmogorov–Arnold Networks (PIKANs). These models are particularly well-suited for solving inverse problems, where the objective is to infer unknown quantities—such as parameters ($p$), boundary conditions, or unobserved inputs—by leveraging observed data alongside prior physical knowledge.
@@
In the context of PINs for gray-box discovery, the goal is to estimate the missing component of the equations—the unknown dynamics of the model ($f(t)$)—in a physical system described by a PDE or ODE. A typical formulation involves a network that aims to minimize a loss function, which is based on the physical model and the available data. Concretely, suppose we have a differential operator
@@
In this study, we compare the performance of two different representation networks for $\hat{u}(t;\theta)$ and $f(t)$: the conventional multilayer perceptron (MLP) and the Kolmogorov–Arnold Network (KAN). In the following sections, we provide details on these two architectures and discuss their respective advantages for capturing the missing dynamics in gray-box discovery scenarios. The detailed architectures of PINNs and PIKANs are shown in Fig. 1.
Figure 1: Physics-Informed Networks.
@@ 2.1 Kolmogorov-Arnold Networks Representation Model
Motivated by the Kolmogorov–Arnold representation theorem, Kolmogorov–Arnold Networks (KANs) have been introduced as a structured alternative to conventional Multi-Layer Perceptrons (MLPs) [51]. This theorem states that any continuous function defined on a compact $d$-dimensional domain, $f:[0,1]^{d}\rightarrow\mathbb{R}$, can be decomposed into a finite combination of nested univariate continuous functions:
@@
Physics-Informed KANs (PIKANs) for forward modeling were systematically explored in [52, 50], where each layer of the network is defined by:
@@
Among various architectural variants, [52, 53] introduced cPIKANs, which use Chebyshev polynomials to parameterize the inner and outer functions. A typical inner function in this setting is given by:
@@
where $\zeta\in[-1,1]$ is the input, $T_{n}$ is the $n$-th Chebyshev polynomial defined recursively as $T_{0}(\zeta)=1,T_{1}(\zeta)=\zeta,T_{n}(\zeta)=2\zeta T_{n-1}(\zeta)-T_{n-2}(\zeta)$, and $\theta=\{c_{n}\}$ are learnable coefficients. This spectral representation has been shown to improve robustness and data efficiency.
@@
$f(\bm{x})\approx W\cdot\sigma\left(\sum_{i_{L}=1}^{m_{L}}\sum_{n_{L}=1}^{d}c_{n_{L}}T_{n_{L}}\left(\sigma\left(\cdots\sum_{i_{1}=1}^{m_{1}}\sum_{n_{1}=1}^{d}c_{n_{1}}T_{n_{1}}\left(\sigma\left(\sigma\left(\sum_{i_{0}=1}^{m_{0}}\sum_{n_{0}=1}^{d}c_{n_{0}}T_{n_{0}}(\sigma(x))\right)\right)\right)\right)\right)\cdots\right)$ (4)
In our architecture, $d$ denotes the degree of the polynomials. The set $\{m_{i}\}_{i=0}^{L}$ represents the number of nodes in the $i^{\text{th}}$ layer, where $L$ is the total number of layers. The trainable parameters of our network, $\theta$, include both the Chebyshev coefficients $\{c_{j}\}_{j=0}^{L}$ and the final weight matrix $W$. The matrix $W$ has dimensions equal to the number of nodes in the last hidden layer multiplied by the number of outputs.
In our network, applying a second $\tanh$—i.e., computing $\tanh(\tanh(x))$—further contracts the activation range from $[-1,1]$ to approximately $[-0.76,0.76]$ (since $\tanh(1)\approx 0.7616$ and $\tanh(-1)\approx-0.7616$). The $\tanh$ function is defined as
$\tanh(x)=\frac{e^{x}-e^{-x}}{e^{x}+e^{-x}},$
and its derivative is given by
$\tanh^{\prime}(x)=1-\tanh(x)^{2}.$
This additional non-linear mapping, even on inputs already confined to $[-1,1]$, refines feature representations and stabilizes learning by further squashing extreme values, which smooths the gradient flow (as gradients near the saturation limits decrease) and provides implicit regularization. This controlled contraction of the activation range ensures that the inputs to subsequent layers—such as those employing Chebyshev polynomial expansions—remain within an optimal domain, thereby enhancing numerical stability and improving the model’s approximation accuracy. Additionally, by applying $\tanh$ at the last hidden layer, all nodes output are mapped to $[-1,1]$, and the scaling multiplication by $W$ is then used to adjust these values to the required range for our problem.
@@
2.2 Multi-Layer Perceptron Representaion Model

### v2

@@ Abstract
Physics-Informed Kolmogorov–Arnold Networks (PIKANs) have been gaining attention as an effective counterpart to the original multilayer perceptron-based Physics-Informed Neural Networks (PINNs). Both representation models can address inverse problems and facilitate gray-box system identification. However, a comprehensive understanding of their performance in terms of accuracy and speed remains underexplored. In particular, we introduce a modified PIKAN architecture, tanh-cPIKAN, which is based on Chebyshev polynomials for parametrization of the univariate functions with an extra nonlinearity for enhanced performance. We then present a systematic investigation of how the choices of optimizer, representation, and training configuration influence the performance of PINNs and PIKANs in the context of systems pharmacology modeling. We benchmark a wide range of optimizers using simple but representative pharmacokinetic and pharmacodynamic models. We use the new Optax library [deepmind2020jax] but also a new class of self-scaled optimizers developed in Optimistix library [optimistix] to identify the most effective combinations for learning gray-boxes under ill-posed, non-unique, and data-sparse conditions. We examine the influence of model architecture (MLP vs. KAN), numerical precision (single vs. double), the need for warm-up phases for second-order methods, and sensitivity to the initial learning rate. We also assess the optimizer scalability for larger models and analyze the trade-offs introduced by JAX in terms of computational efficiency and numerical accuracy. Using two representative system pharmacology examples, a pharmacokinetics model and a chemotherapy drug response model, we offer practical guidance on selecting optimizers and representation models/architectures for robust and efficient gray-box discovery. Our findings provide actionable insights for improving the training of physics-informed networks in systems pharmacology, systems biology, and beyond.
@@
PINNs, PIKANs, Kolmogorov-Arnold Networks, Systems Biology, Systems Pharmacology, Pharmacometrics
@@ 1 Introduction
Inverse problems are omni-present across scientific domains, enabling the inference of hidden parameters and dynamic system behaviors from empirical observations. Such problems are especially significant in biology, physics, and pharmacology, where direct measurements are often impractical, and systems exhibit stiff, nonlinear dynamics. Traditional methods such as Sparse Identification of Nonlinear Dynamics (SINDy) [brunton2016discovering], which build on sparse regression [tibshirani1996regression] and compressed sensing [donoho2006compressed], have demonstrated success in extracting interpretable equations from data. However, their reliance on predefined function libraries limits their expressiveness. Recent extensions like ADAM-SINDy [adamSINDy2024] address this by jointly optimizing coefficients and nonlinear parameters via first-order methods such as Adam, thereby enhancing robustness.
Other symbolic regression methods—including AI-Feynman [udrescu2020ai], Feyn [brolos2021approach, christensen2022identifying], and AI-Descartes [cornelio2023combining]—learn compact equations directly from data without assuming specific model structures. While effective in low-data regimes, they often struggle when systems involve differential or integral operators. To overcome these challenges, hybrid modeling approaches that integrate data-driven and physics-based priors have emerged. Notably, physics-informed neural networks (PINNs) [raissi2019physics] and their variants (e.g., X-PINNs [jagtap2021extended]) enforce known physical laws during training and have been combined with symbolic methods for gray-box system discovery [kiyani2023framework]. Here by gray-box modeling we refer to hybrid frameworks that combine known mechanistic equations with neural components that learn unknown or uncertain dynamics from data; importantly, the missing part of the governing equations cannot be discovered directly from data, since the data describe the solution of the differential system rather than its underlying operators, and must therefore be inferred indirectly through physics-constrained learning. Additional contributions include random projection neural networks [de2018greedy, de2021interaction] and symbolic-layer least-squares models [kokturk2021symbolic], which improve scalability and interpretability.
Early gray-box modeling using neural networks [rico1994continuous] has since evolved into frameworks for applications including phase-field models, optogenetics, and metabolic systems [kemeth2023black, lovelett2019partial]. In biology, parameter inference from noisy and sparse data has motivated the development of systems biology-informed neural models [yazdani2020systems, daneker2023systems]. AI-Aristotle [Ahmadi2024] exemplifies this trend by integrating domain-decomposed PINNs with symbolic regression [raissi2024physics] to perform gray-box modeling under data constraints. Similarly, CMINNs [daryakenari2025cminns] combine (fractional) PINNs with compartmental models to estimate time-varying PK/PD parameters, enabling the modeling of non-standard drug behaviors such as anomalous diffusion, delayed response, and resistance. These advances offer flexible representations for inverse problems in biomedical systems where parameter constancy assumptions may fall short.
Despite these developments, training PINNs and their variants remains a key bottleneck. Their composite loss functions—balancing physics residuals and data misfit—result in non-convex optimization landscapes prone to local minima and slow convergence. This is further exacerbated in systems pharmacology, where datasets are sparse, unbalanced, and noisy, and the parameters of interest are often time-dependent or embedded within nonlinear dynamics. Consequently, optimization algorithms play a crucial role in the performance and reliability of these models.
Current training methods rely heavily on first-order optimizers such as stochastic gradient descent and Adam [34, 35], which, while scalable, often plateau prematurely. High-order methods like Newton’s method offer faster convergence by leveraging curvature but are constrained by memory and computation costs due to the Hessian. To mitigate this, quasi-Newton methods such as BFGS and L-BFGS approximate the inverse Hessian using rank-one or rank-two updates [41, 42, 43], and have been adapted to PINN training with some success [elham]. Recent studies have also explored adaptive loss weighting, warm-up strategies, exponential feature scaling, and sequential training as means to improve convergence and stability [Ahmadi2024, URBAN2025113656].
In this work, we systematically examine the optimization landscape of gray-box inverse problems in systems biology. Specifically, we investigate the impact of various optimizers—including adaptive first-order methods with learning rate schedulers and second-order methods like BFGS with backtracking and trust-region line search—on training dynamics and solution quality. Additionally, we assess the role of numerical precision (float32 vs. float64) in convergence behavior and final model accuracy. Our findings provide actionable insights for developing stable and interpretable physics-informed models in pharmacology and broader biological contexts.
@@
The paper is organized as follows. In Section 2, we present the methodology, detailing the architectures of PINNs and the proposed tanh-cPIKANs, along with their respective formulations and training strategies. In Section 3, we introduce representative pharmacokinetics and pharmacodynamics models used for evaluation, both of which represent gray-box inverse problems characterized by data sparsity, non-uniqueness, and ill-posedness. In Section 4, we provide a systematic investigation into representation and optimization choices for gray-box discovery. We compare the performance of MLP-based PINNs and KAN-based architectures, explore the influence of numerical precision (single vs. double) on convergence and stability, and evaluate the effectiveness of various optimizers and learning rate schedulers. Moreover, we assess the computational cost and accuracy trade-offs associated with using the JAX framework. Finally, in Section 5 we summarize the key findings, emphasizing the advantages of the proposed tanh-cPIKAN architecture to improve training stability and accuracy for gray-box system identification.
@@ 2 Methodology
Physics-Informed Networks (PINs) are models that embed physical laws—typically represented by ordinary or partial differential equations (ODEs/PDEs)—directly into the training process of function approximators such as neural networks or Kolmogorov–Arnold Networks. This is accomplished through automatic differentiation, which enables the computation of derivatives required to enforce consistency with the governing equations. The term “PINs” broadly encompasses both Physics-Informed Neural Networks (PINNs) and Physics-Informed Kolmogorov–Arnold Networks (PIKANs). These models are particularly well-suited for solving inverse problems, where the objective is to infer unknown quantities—such as parameters, boundary conditions, or unobserved inputs—by leveraging observed data alongside prior physical knowledge.
@@
In the context of PINs for gray-box discovery, the goal is to estimate the missing component of the equations—the unknown dynamics of the model (f(t)f(t))—in a physical system described by a PDE or ODE. A typical formulation involves a network that aims to minimize a loss function, which is based on the physical model and the available data. Concretely, suppose we have a differential operator
@@
In this study, we compare the performance of two different representation networks for $\hat{u}(t;\theta)$ and $f(t)$: the conventional multilayer perceptron (MLP) and the Kolmogorov–Arnold Network (KAN). In the following sections, we provide details on these two architectures and discuss their respective advantages for capturing the missing dynamics in gray-box discovery scenarios. Schematics of the architectures of PINNs and PIKANs are shown in Fig. 1.
Figure 1: PINs: Physics-Informed Networks. Here $u$ and $u_{0}$ are observables at $t>0$ and $t=0$, respectively. $\hat{u}$ and $\hat{u}_{0}$ are inferred field from PINNs or PIKANs at t>0 and $t=0$, respectively.
@@ 2.1 Kolmogorov-Arnold Networks Representation Model
Motivated by the Kolmogorov–Arnold representation theorem, Kolmogorov–Arnold Networks (KANs) have been introduced as a structured alternative to conventional Multi-Layer Perceptrons (MLPs) [liu2024kan]. This theorem states that any continuous function defined on a compact $d$-dimensional domain, $f:[0,1]^{d}\rightarrow\mathbb{R}$, can be decomposed into a finite combination of nested univariate continuous functions:
@@
Physics-Informed KANs (PIKANs) for forward modeling were systematically explored in [shukla2024comprehensive, review1, ahmadi202new], where each layer of the network is defined by:
@@
Among various architectural variants, [shukla2024comprehensive, ss2024chebyshev] introduced cPIKANs, which use Chebyshev polynomials to parameterize the inner and outer functions. A typical inner function in this setting is given by:
@@
where $\zeta\in[-1,1]$ is the input, $T_{n}$ is the $n$-th Chebyshev polynomial defined recursively as $T_{0}(\zeta)=1,T_{1}(\zeta)=\zeta,T_{n}(\zeta)=2\zeta T_{n-1}(\zeta)-T_{n-2}(\zeta)$, and $\theta=\{c_{n}\}$ are learnable coefficients. This spectral representation has been shown to improve robustness and computational efficiency. In the context of physics-informed training, our experiments demonstrate that Chebyshev polynomial bases outperform Jacobi variants: despite the improved expressivity reported in Ref. [AFZAL] for broader machine-learning tasks (rather than PINNs), our parameter-matched results in  A.1 show that Jacobi-based KANs incur substantially longer runtime—due to their parameter-dependent recurrence coefficients, additional normalization factors, and larger computational graphs that hinder kernel fusion—while offering no accuracy improvement over the Chebyshev-based PIKANs.
@@
$f(\bm{x})\approx W\cdot\sigma\left(\sum_{i_{L}=1}^{H_{L}}\sum_{n_{L}=1}^{D}c_{n_{L}}T_{n_{L}}\left(\sigma\left(\cdots\sum_{i_{1}=1}^{H_{1}}\sum_{n_{1}=1}^{D}c_{n_{1}}T_{n_{1}}\left(\sigma\left(\sigma\left(\sum_{i_{0}=1}^{H_{0}}\sum_{n_{0}=1}^{D}c_{n_{0}}T_{n_{0}}(\sigma(x))\right)\right)\right)\right)\right)\cdots\right)$ (4)
In our architecture, $D$ denotes the degree of the polynomials. The set $\{H_{i}\}_{i=0}^{L}$ represents the number of nodes in the $i^{\text{th}}$ layer, where $L$ is the total number of layers. The trainable parameters of our network, $\theta$, include both the Chebyshev coefficients $\{c_{j}\}_{j=0}^{L}$ and the final weight matrix $W$. The matrix $W$ has dimensions equal to the number of nodes in the last hidden layer multiplied by the number of outputs.
In our network, applying a second $\tanh$—i.e., computing $\tanh(\tanh(x))$—further contracts the activation range from [1,1][-1,1] to approximately [0.76,0.76][-0.76,0.76] (since $\tanh(1)\approx 0.7616$ and $\tanh(-1)\approx-0.7616$). This additional non-linear mapping, even on inputs already confined to [1,1][-1,1], refines feature representations and stabilizes learning by further squashing extreme values, which smooths the gradient flow (as gradients near the saturation limits decrease) and provides implicit regularization. This controlled contraction of the activation range ensures that the inputs to subsequent layers—such as those employing Chebyshev polynomial expansions—remain within an optimal domain, thereby enhancing numerical stability and improving the model’s approximation accuracy. Additionally, by applying $\tanh$ at the last hidden layer, all nodes output are mapped to [1,1][-1,1], and the scaling multiplication by $W$ is then used to adjust these values to the required range for our problem.
@@
2.2 Multi-Layer Perceptron Representation Model